# Evaluating the effects of Data Sparsity on the Link-level Bicycling Volume Estimation: A Graph Convolutional Neural Network Approach

### v1

@@ 4.5 Sparsity Simulation
To systematically assess the impact of data sparsity on the performance of our GCN model, we introduced varying levels of sparsity into the dataset. Sparsity is simulated by selectively masking a percentage of nodes in the training data i.e. intentionally removing a certain proportion of the labelled nodes with Strava AADB counts from model training. This is performed to mimic the real-world scenario where data collection might be incomplete and sparse.
The removal process is performed randomly to ensure that the sparsity introduced is unbiased and reflective of potential missing data scenarios in bicycling networks. We have applied different sparsity levels - 20%, 50%, 60%, 70%, 80%, 90% and 98.9%, to create datasets with progressively fewer data points, ultimately reaching a condition that mirrors the ground truth scenario in the City of Melbourne. For each sparsity level, we recalculate the connectivity and structure of the graph to ensure that the remaining data still forms a coherent network. These sparse dataset is then used to train and evaluate the GCN model, allowing us to observe how the model’s predictive performance is affected with increasing sparsity.
@@ 5.3 Effects of data sparsity on the models performance
This section presents a detailed evaluation of how varying levels of data sparsity affect the performance of the Graph Convolutional Network (GCN) compared to traditional machine learning models LR, SVM, RF — in estimating link-level Annual Average Daily Bicycle (AADB) counts. The results of our experiments under different sparsity levels, from 0% to 98.9%, are summarized in Table 5, and the corresponding performance metrics are visualized in Figure 2.
@@
Extreme Sparsity Levels (90% - 98.9%): At 90% sparsity, with 1,274 nodes for training, the GCN performance starts to significantly degrade, with an RMSE of 32.507 suggesting that the effectiveness of GCN diminishes when data becomes excessively sparse. RF exhibits a more gradual decline, achieving an RMSE of 26.707. The reduced complexity of RF allows it to handle extreme sparsity more reliably, though with reduced accuracy compared to GCN at lower sparsity levels. At extreme sparsity levels (98.9%), the labeled nodes drop to only 141 which is similar to the real-world scenario of count locations in the City of Melbourne. At this level, all models suffer from high RMSE, MAE, and MAPE values, but GCN’s performance falls more drastically, with RMSE reaching 44.838. The RF model, though still not ideal, becomes relatively more reliable under these conditions, with an RMSE of 32.392. The inability of GCN to rely on spatial aggregation due to a lack of sufficient labeled nodes results in its significant decline.
@@
6 Conclusion
This study addresses the unique challenges of link-level bicycling volume estimation by leveraging a Graph Convolutional Network (GCN) architecture specifically designed for the geographical area of the City of Melbourne. While traditional link-level volume estimation models have demonstrated effectiveness in motorized traffic applications, they fall short in the bicycling context due to the sparse nature of the available data and the complex mobility patterns unique to bicycling. To the best of our knowledge, this is the first study to address the data sparsity challenge and utilize a Graph Convolutional Network (GCN) architecture specifically to model link-level bicycling volumes in a real-world urban context, thereby filling a significant gap in the existing literature.
By integrating OpenStreetMap (OSM) network data with Strava Metro’s comprehensive cycling activity data, we simulate varying levels of data sparsity and evaluate the robustness of the GCN model under these conditions. Our findings reveal that the GCN outperforms traditional machine learning models particularly in scenarios with low to moderate data sparsity. Specifically, the optimal GCN configuration achieved the lowest Root Mean Squared Error (RMSE) across different sparsity levels, underscoring its superior ability to capture spatial dependencies inherent in bicycling networks.
@@ 6 Conclusion
Future research should focus on enhancing the robustness of GCNs in extremely sparse data environments. Potential avenues include the development of hybrid models that combine GCNs with other machine learning techniques, the incorporation of auxiliary data sources to enrich the feature set, and the exploration of advanced data augmentation methods to compensate for missing data. Additionally, expanding the study to other urban contexts can validate the generalizability of the proposed approach and further refine the model’s applicability across diverse bicycling networks.
@@ A.2 Training and Validation Loss Curve for Model training at different sparsity levels
Figure 4 presents the training and validation loss curves for the GCN model (configuration G) across various levels of data sparsity, ranging from 0% to 98.89%. These plots are essential for understanding how the GCN model handles increasing data sparsity, which mirrors real-world scenarios where count data is extremely sparse.
As data sparsity increases, the model’s learning process becomes more challenging. The earlier onset of early stopping, as seen at higher sparsity levels like 90% and 98.89%, suggests that the model struggles to find sufficient signal from the limited data available. The validation loss curves also show increased volatility as sparsity increases, indicating a difficulty in maintaining generalization.
@@
Figure 4: Loss curves for GCN under different levels of sparsity

### v2

@@ 4.5 Sparsity Simulation
To systematically assess the impact of data sparsity on the performance of our GCN model, we introduced varying levels of sparsity into the dataset as shown in Figure 4. Sparsity is simulated by selectively masking a percentage of nodes in the training data i.e. intentionally removing a certain proportion of the labelled nodes with Strava AADB counts from model training. This is performed to mimic the real-world scenario where data collection might be incomplete and sparse.
The removal process is performed randomly to ensure that the sparsity introduced is unbiased and reflective of potential missing data scenarios in bicycling networks. We have applied different sparsity levels - 20%, 50%, 60%, 70%, 80%, 90% and 99%, to create datasets with progressively fewer data points, ultimately reaching a condition that mirrors the ground truth scenario in the City of Melbourne. For each sparsity level, we recalculate the connectivity and structure of the graph to ensure that the remaining data still forms a coherent network. These sparse dataset is then used to train and evaluate the GCN model, allowing us to observe how the model’s predictive performance is affected with increasing sparsity.
@@
(a) Bicycle Network in City of Melbourne
(b) Sparsity - 0%
(c) Sparsity - 20%
(d) Sparsity - 50%
(e) Sparsity - 60%
(f) Sparsity - 70%
(g) Sparsity - 80%
(h) Sparsity - 90%
(i) Sparsity - 99%
Figure 4: Visualization of sparsity simulation for the bicycle network in the City of Melbourne. (a) Full bicycle network representation. (b–i) Graphs demonstrating different sparsity levels: 0%, 20%, 50%, 60%, 70%, 80%, 90%, and 99%. Red indicates training data, yellow indicates validation data, and green indicates testing data. As sparsity increases, fewer nodes are retained for training, simulating real-world scenarios of incomplete data. This visualization highlights the systematic reduction in training nodes.
@@ 5.3 Effects of data sparsity on the models performance
This section presents a detailed evaluation of how varying levels of data sparsity affect the performance of the Graph Convolutional Network (GCN) compared to traditional machine learning models LR, SVM, RF — in estimating link-level Annual Average Daily Bicycle (AADB) counts. The results of our experiments under different sparsity levels, from 0% to 99%, are summarized in Table 5, and the corresponding performance metrics are visualized in Figure 5.
@@
Extreme Sparsity Levels (90% - 99%): At 90% sparsity, with 1,274 nodes for training, the GCN performance starts to significantly degrade, with an RMSE of 32.507 suggesting that the effectiveness of GCN diminishes when data becomes excessively sparse. RF exhibits a more gradual decline, achieving an RMSE of 26.707. The reduced complexity of RF allows it to handle extreme sparsity more reliably, though with reduced accuracy compared to GCN at lower sparsity levels. At extreme sparsity levels (99%), the labeled nodes drop to only 127 which is similar to the real-world scenario of count locations in the City of Melbourne. At this level, all models suffer from high RMSE, MAE, and MAPE values, but GCN’s performance falls more drastically, with RMSE reaching 44.838. The RF model, though still not ideal, becomes relatively more reliable under these conditions, with an RMSE of 32.392. The inability of GCN to rely on spatial aggregation due to a lack of sufficient labeled nodes results in its significant decline.
6 Conclusion and Future Research Directions
This study addresses the unique challenges of link-level bicycling volume estimation by leveraging a Graph Convolutional Network (GCN) architecture specifically designed for the geographical area of the City of Melbourne. While traditional link-level volume estimation models have demonstrated effectiveness in motorized traffic applications, they fall short in the bicycling context due to the sparse nature of the available data and the complex mobility patterns unique to bicycling. To the best of our knowledge, this is the first study to evaluate the effects of data sparsity and utilize a Graph Convolutional Network (GCN) architecture specifically to model link-level bicycling volumes in a real-world urban context, thereby filling a significant gap in the existing literature.
By integrating OpenStreetMap (OSM) network data with Strava Metro’s bicycling activity data, we simulate varying levels of data sparsity and evaluate the robustness of the GCN model under these conditions. Our findings reveal that the GCN outperforms traditional machine learning models particularly in scenarios with low to moderate data sparsity. Specifically, the optimal GCN configuration achieved the lowest Root Mean Squared Error (RMSE) across different sparsity levels, underscoring its superior ability to capture spatial dependencies inherent in bicycling networks.
@@ 6 Conclusion and Future Research Directions
Future research should prioritize enhancing the robustness of Graph Convolutional Networks (GCNs) in environments characterized by extreme data sparsity. One promising approach is the development of hybrid models that combine GCNs with other machine learning techniques, leveraging their complementary strengths to improve performance under high sparsity conditions. Additionally, the incorporation of auxiliary data sources—such as real-time traffic sensors, environmental factors, and demographic information—can significantly enrich the feature set, compensating for missing data and enhancing the model’s predictive capabilities.
Exploring advanced data augmentation methods is another critical avenue, as sophisticated techniques can artificially increase data density and provide more diverse training samples, thereby improving model generalization. Moreover, expanding the study to ensure generalizability across diverse urban contexts is essential. Applying the proposed GCN approach to other cities with varied bicycling networks will validate its adaptability and effectiveness in different settings, ensuring broader applicability.
Addressing and mitigating data biases inherent in crowd-sourced data, such as those from Strava, is also crucial. Developing methods to correct these biases will ensure more accurate and representative model predictions, enhancing the reliability of bicycling volume estimates. Furthermore, integrating temporal dynamics into the model by incorporating temporal layers or time-series data will enhance its ability to capture temporal patterns, allowing for more dynamic and responsive predictions that reflect fluctuating bicycling demand over time.
Lastly, extending the model to support real-time prediction capabilities represents a significant advancement. Real-time bicycling volume predictions can be instrumental for immediate infrastructure adjustments and traffic management, providing urban planners and policymakers with timely insights to promote sustainable transportation. By pursuing these research directions, future studies can further refine and extend the applicability of GCNs in bicycling volume estimation, making them even more valuable for urban mobility planning and sustainable transportation initiatives.
@@
Acknowledgements
The CYCLED (CitY-wide biCycLing Exposure modelling) Study is funded by an Australian Research Council Discovery Project (DP210102089). Mohit Gupta’s PhD scholarship and Debjit Bhowmick were supported by the Australian Research Council Discovery Project (DP210102089). Ben Beck was supported by an Australian Research Council Future Fellowship (FT210100183). This work includes aggregated and de-identified data from Strava Metro (Strava Metro contributors, , 2022).
@@ A.2 Training and Validation Loss Curve for Model training at different sparsity levels
Figure 7: Loss curves for GCN under different levels of sparsity
Figure 7 presents the training and validation loss curves for the GCN model (configuration G) across various levels of data sparsity, ranging from 0% to 99%. These plots are essential for understanding how the GCN model handles increasing data sparsity, which mirrors real-world scenarios where count data is extremely sparse.
As data sparsity increases, the model’s learning process becomes more challenging. The earlier onset of early stopping, as seen at higher sparsity levels like 90% and 99%, suggests that the model struggles to find sufficient signal from the limited data available. The validation loss curves also show increased volatility as sparsity increases, indicating a difficulty in maintaining generalization.